What are the steps to finding 4(log9 of X)^2 - 2(log9 of X^2)=3?
Answer by stanbon(26268)

What are the steps to finding
4(log9 of X)^2 - 2(log9 of X^2)=3?
------------------
It is a quadratic with w = (log9 (x))
-------------
Rewrite as:
4w^2 - 2w - 3 = 0
--
w = [2 +- sqrt(4 -4*4*-3)]/8
---
w = [2 +- sqrt(52)]/8
---
w = [2 +- 2sqrt(13)]/8
---
w = [(1/4) + (1/4)sqrt(13)] or w = [(1/4) - (1/4)sqrt(13)]
----------------
Now solve for "x"
log9 (x) = [(1/4) + (1/4)sqrt(13)] or log9 (x) = [(1/4) - (1/4)sqrt(13)]
---
x = 9^[(1/4) + (1/4)sqrt(13)] or x = 9^[(1/4) - (1/4)sqrt(13)]
---
x = 12.55.. or x = 0.239...
=================================
Cheers,
Stan H.